Hello, could you please show the steps and answer for s in the equation
3s + 5 = -21
Thank you,Peggy

Answer by farohw(175)


Hello,

3s + 5 = -21

3s = -21 - 5

3s = -26

Divide both sides by 3 to isolate s:
3s/3 = -26/3

s = -26/3

You can verify this by putting s = -26/3 back into the original equation:
3s + 5 = -21 --> 3(-26/3) + 5 = -21

-26 + 5 = -21 --> -21 = -21